OpenStudy (anonymous):

tan^2(x)=1 − cos(2x)/1 + cos(2x)

OpenStudy (anonymous):

tan^2x= 1- cos2x/1+cos2x tan^2x= sin^2x/cos^2x sin^2x=1-cos2x/2 cos^2x= 1+ cos2x/2 tan^2x=1-cos2x/1+cos2x